# DRAGON: Distributional Rewards Optimize Diffusion Generative Models

### v1

@@ Abstract
We present Distributional RewArds for Generative OptimizatioN (DRAGON), a versatile framework for fine-tuning media generation models towards a desired outcome. Compared with traditional reinforcement learning with human feedback (RLHF) or pairwise preference approaches such as direct preference optimization (DPO), DRAGON is more flexible. It can optimize reward functions that evaluate either individual examples or distributions of them, making it compatible with a broad spectrum of instance-wise, instance-to-distribution, and distribution-to-distribution rewards. Leveraging this versatility, we construct novel reward functions by selecting an encoder and a set of reference examples to create an exemplar distribution. When cross-modality encoders such as CLAP are used, the reference examples may be of a different modality (e.g., text versus audio). Then, DRAGON gathers online and on-policy generations, scores them to construct a positive demonstration set and a negative set, and leverages the contrast between the two sets to maximize the reward. For evaluation, we fine-tune an audio-domain text-to-music diffusion model with 20 different reward functions, including a custom music aesthetics model, CLAP score, Vendi diversity, and Fréchet audio distance (FAD). We further compare instance-wise (per-song) and full-dataset FAD settings while ablating multiple FAD encoders and reference sets. Over all 20 target rewards, DRAGON achieves an $81.45\%$ average win rate. Moreover, reward functions based on exemplar sets indeed enhance generations and are comparable to model-based rewards. With an appropriate exemplar set, DRAGON achieves a $60.95\%$ human-voted music quality win rate without training on human preference annotations. As such, DRAGON exhibits a new approach to designing and optimizing reward functions for improving human-perceived quality. Example generations can be found at https://ml-dragon.github.io/web.
@@ 1 Introduction
22footnotetext: Work done as an intern at Adobe Research.
@@
Recent advances in diffusion models have transformed content generation across media domains, establishing new standards for generating high-quality images, video, and audio (Rombach et al., 2022; Ho et al., 2022; Liu et al., 2023; Ghosal et al., 2023). While these models achieve impressive results through sophisticated training schemes, their optimization process typically focuses on metrics that may not align with downstream objectives or human preferences. This misalignment creates a fundamental challenge: how can we effectively steer these models toward desired output distributions or optimize them for specific performance metrics?
@@
A prominent approach to address this challenge has been fine-tuning using instance-level feedback. Methods such as reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Williams, 1992; Schulman et al., 2017) and related methods like Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) leverage pre-trained reward models or large-scale, pairwise preference data collected offline to guide the optimization process. While effective, these approaches face several key challenges. First, media like audio, music, and video are multi-modal and highly perceptual, making it not only hard and expensive to create reliable preference pairs (e.g. music (Cideron et al., 2024), in-the-wild audio (Liao et al., 2024)), but also challenging to adopt criteria-based reward signals, as recently popularized in language model training (Guo et al., 2025). Second, preference-based training methods are constrained by the implicit reward functions hidden in their training data, making it difficult to adapt these approaches to new objectives or target distributions without collecting new preference data. Third, these approaches are not fully connected to the generative model evaluation metrics that often measure distributional properties like Fréchet embedding distance, diversity, and coverage.
@@
To address these limitations, we introduce Distributional RewArds for Generative OptimizatioN (DRA-GON), a versatile framework for fine-tuning generative models towards a desired outcome or target distribution. DRAGON offers an alternative to existing reinforcement learning (RL) methods or pair-wise preference approaches for optimizing a broad spectrum of rewards, including instance-wise, instance-to-distribution, and distribution-to-distribution signals. As shown in Figure 1(a), the key components of DRAGON are: 1) a pre-trained embedding extractor and a set of (possibly cross-modal) reference examples, 2) a reward-based scoring mechanism that creates positive and negative sets of on-policy online generations, and 3) an optimization process that leverages the contrast between the positive and negative sets. To our knowledge, DRAGON is the first practical algorithm that reliably optimizes our entire taxonomy of reward functions for generative models. We believe the ability to handle distribution-to-distribution rewards is particularly valuable since we can directly optimize generation quality metrics. Such reward functions are ubiquitous in standard generative model evaluation metrics (Fréchet embedding distance (Heusel et al., 2017), Kullback–Leibler (KL) divergence (Kullback & Leibler, 1951), Inception score (Salimans et al., 2016)). Moreover, leveraging DRAGON’s unique versatility, we can construct new reward functions by simply collecting a set of ground-truth examples without human preference, drastically reducing the effort to construct reward signals.
@@
We demonstrate DRAGON’s effectiveness through comprehensive experiments with text-to-music diffusion transformers. Our evaluation incorporates multiple common music generation metrics: audio-text alignment (CLAP score) (Elizalde et al., 2023; Wu* et al., 2023), Fréchet audio distance (FAD) (Kilgour et al., 2019) evaluated across diverse reference sets and embedding models, Vendi score (Friedman & Dieng, 2023) for reference-free output diversity, and a custom-built human preference model for reference-free aesthetics scoring, totaling 20 reward functions. As shown in Figure 1(b), DRAGON consistently improves over baselines across these diverse reward functions, achieving an average of 81.45% target reward win rate while generalizing the improvement across evaluation metrics. Via human listening test, we show that DRAGON can achieve a 61.2% win rate in human-perceived music quality without training on human preference data.
@@
(a) Overall diagram of DRAGON, a versatile on-policy learning framework for media generation models that can optimize various types of reward functions.
@@
(b) DRAGON significantly improves a full suite of rewards. Each vertex of the plot considers a reward metric and reports the win rate of the DRAGON model optimized for the metric.
@@
  • We propose a human aesthetics preference model for AI-generated music and find that DRAGON can
    leverage it to improve human-perceived music quality with a small set of (1,676) human-rated clips.
@@ 2.1 Music Generation
Music generation has seen significant advances via auto-regressive token-based models (Vaswani et al., 2017; Zeghidour et al., 2021; Agostinelli et al., 2023; Borsos et al., 2023; Copet et al., 2023) and (latent) diffusion or flow-matching models (Novack et al., 2024a; Huang et al., 2023; Evans et al., 2024b; Wu et al., 2024; Evans et al., 2024a). Among auto-regressive approaches, MusicLM (Agostinelli et al., 2023) and MusicGen (Copet et al., 2023) are notable, with the former extended with RL via MusicRL (Cideron et al., 2024). Latent diffusion models like Stable Audio (Evans et al., 2024b; a) additionally show high-quality results and potential for ultra-fast generation (Bai et al., 2024; Novack et al., 2024b; 2025). However, reward optimization for diffusion models has proven more challenging than for auto-regressive ones (Wallace et al., 2024).
@@ 2.2 Diffusion Models
Diffusion models have emerged as a powerful paradigm able to generate high-quality and diverse samples (Sohl-Dickstein et al., 2015; Ho et al., 2020) via iterative de-noising. Such models commonly consist of a de-noising network $f_{\theta}$ that inputs noisy input $x_{t}$, diffusion time step $t$, and condition $c$ (e.g. text), and can either be discrete-time (Ho et al., 2020) or continuous-time score models (Song et al., 2021b; Karras et al., 2022). During training, given a demonstrative example $x_{0}$, we add Gaussian noise to form the noisy example $x_{t}$, and train $f_{\theta}$ to undo the noising. At inference time, generation begins with pure noise which is repeatedly de-noised. That is, $f_{\theta}$ is applied repeatedly to gradually de-noise, forming a trajectory along the time step $t$ and noise level. Latent diffusion models (Rombach et al., 2022), where the above diffusion process takes place in a latent embedding space, have been leveraged across nearly all perceptual generation domains, including creating images (Rombach et al., 2022; Karras et al., 2022), video (Ho et al., 2022), speech (Popov et al., 2021; Eskimez et al., 2024), in-the-wild audio (Liu et al., 2023; Bai et al., 2024), and music (Huang et al., 2023; Forsgren & Martiros, 2022). Recent advances have shown that transformer-based architectures are advantageous, leading to diffusion transformers (DiT) (Peebles & Xie, 2023).
@@ 2.3 Reward Optimization for Diffusion Models
Diffusion models present unique challenges for reward-based optimization due to their iterative de-noising process. Existing work tackled this by formulating the diffusion process as a Markov Decision Process, with the reward signal assigned either exclusively to the final time step (Fan et al., 2024) or to all time steps (Black et al., 2024). More recent works explored the alternative of implicit reward optimization via learning from contrastive demonstrations. Diffusion-DPO (Wallace et al., 2024) and MaPO (Hong et al., 2024) optimize rewards defined with explicit binary preferences between paired samples (e.g., media contents with the same captions). Diffusion-KTO (Li et al., 2024) offers flexibility by being compatible with unpaired collections of preferred and non-preferred samples. Concurrent to our work, TangoFlux (Hung et al., 2024) represents an application of a CLAP score reward in diffusion modeling, proposing a semi-on-policy approach for in-the-wild audio generation.
@@ 2.4 Human Feedback Datasets and Aesthetics Models
Human-feedback datasets and aesthetics models play a crucial role in guiding the optimization of generative models towards human preferences and reward models. Datasets such as SAC (Pressman et al., 2022), AVA (Murray et al., 2012), LAION-Aesthetics V2 (Schuhmann, 2022), Pick-a-Pic (Kirstain et al., 2023), and RichHF-18K (Liang et al., 2024) have been instrumental in improving the quality and alignment of generated images. These datasets are then used to train aesthetics models, which often map pre-extracted embeddings to a preference score. For audio/music generation, such an approach has been investigated with MusicRL (music) and BATON (in-the-wild audio), but is still rare and limited. In Section 4.1 and Appendix B, we describe the construction of our own aesthetics dataset and reward model for music generation.
@@ 3 Distributional Reward Optimization For Diffusion Models
We propose DRAGON, a reward optimization framework for optimizing diffusion models with a wide variety of reward signals as shown in Figure 1(a). We consider a distributional reward function $r_{\textrm{dist}}:{\mathcal{P}}\to{\mathbb{R}}$ that assigns a reward value to a distribution of generations, where ${\mathcal{P}}$ is the set of all such distributions. We allow $r_{\textrm{dist}}$ to be non-differentiable (e.g., human preference). The outputs of our generative model $f_{\theta}$ form a distribution ${\mathcal{D}}_{\theta}$. When $f_{\theta}$ is a conditional generator, ${\mathcal{D}}_{\theta}$ depends on the distribution of conditioning ${\mathcal{C}}$, although we omit ${\mathcal{C}}$ for notation simplicity. Our goal is to fine-tune a pre-trained model via
@@
The formulation eq. 1 includes the widely studied instance-level reward optimization tasks, such as RLHF. Specifically, we recover instance-level optimization when the distributional reward $r_{\textrm{dist}}({\mathcal{D}}_{\theta})$ is the expectation over an instance-level reward $\mathbb{E}_{X\sim{\mathcal{D}}_{\theta}}[r_{\textrm{instance}}(X)]$. However, our framework extends beyond this special case, which is significant because traditional RL methods such as policy gradients are limited to instance-level rewards, as they cannot distinguish between high and low reward generations without more granular feedback.
To tackle such challenges and optimize $r_{\textrm{dist}}({\mathcal{D}}_{\theta})$, we construct a positive demonstrative distribution ${\mathcal{D}}_{+}$ such that $r_{\textrm{dist}}({\mathcal{D}}_{+})>r_{\textrm{dist}}({\mathcal{D}}_{\theta})$ and a negative demonstrative distribution ${\mathcal{D}}_{-}$. We then optimize model parameters $\theta$ to make ${\mathcal{D}}_{\theta}$ imitate ${\mathcal{D}}_{+}$ and repel ${\mathcal{D}}_{-}$. In the following sections, we address how to: 1) construct ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$, and 2) optimize $\theta$ to make ${\mathcal{D}}_{\theta}$ imitate ${\mathcal{D}}_{+}$ and repel ${\mathcal{D}}_{-}$. Please also find pseudocode in Appendix F.
@@ 3.1 On-Policy Construction of ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$
Existing work often assumes demonstrations ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$ to be provided in advance (offline and off-policy), with ${\mathcal{D}}_{+}$ known to achieve a higher reward than ${\mathcal{D}}_{-}$ (Rafailov et al., 2023; Ethayarajh et al., 2024; Majumder et al., 2024). For instance-level rewards, this can be achieved by splitting a dataset into two halves at a reward threshold, a common RLHF approach. The limitations of off-policy learning necessitate a shift toward on-policy learning for several key reasons. First, for non-instance-level rewards such as FAD, the split becomes less straightforward. Second, large-scale offline data, required for effective off-policy learning, may be unavailable in practice. Third, on-policy optimization disentangles reward from dataset, providing more flexibility in reward choice. Finally, on-policy learning has demonstrated superior effectiveness and robustness because it enables real-time feedback and reduces data-policy mismatch. In the context of generative modeling, Tajwar et al. (2024) showed on-policy data helps language models learn from negative examples, and Hung et al. (2024) showed that even a partially on-policy data collection pipeline improves diffusion models. Hence, we focus on an online and on-policy approach, although offline data can be optionally incorporated into DRAGON with minimal algorithmic changes.
@@
To construct on-policy distributions ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$, we sample from ${\mathcal{D}}_{\theta}$ (which updates throughout training) online and query the distributional reward $r_{\textrm{dist}}$ on the fly. Specifically, before each training step, we collect two batches of observations from ${\mathcal{D}}_{\theta}$ by running full model inference with $f_{\theta}$ and denote them as ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. While the notations ${\mathcal{D}}_{1}$, ${\mathcal{D}}_{2}$, ${\mathcal{D}}_{+}$, and ${\mathcal{D}}_{-}$ technically represent distributions of generations, for simplicity, we also use them to denote the sampled demonstration sets. For the instance-level reward special case, where $r_{\textrm{dist}}=\mathbb{E}_{X\sim{\mathcal{D}}_{\theta}}[r_{\textrm{instance}}(X)]$, the contrastive demonstration sets $({\mathcal{D}}_{+},{\mathcal{D}}_{-})$ can be constructed by taking the better/worse halves of the union ${\mathcal{D}}_{1}\cup{\mathcal{D}}_{2}$. This split can be determined by protocols such as element-level pair-wise comparison (if ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$ consist of paired examples) or comparison with the batch median reward.
@@
Optimizing general rewards like $r_{\textrm{dist}}$ that evaluate distributions of generations is more delicate as we need to disentangle each element’s contribution. To this end, we propose Algorithm 1, a greedy algorithm. We initialize the positive/negative demonstration sets $({\mathcal{D}}_{+}^{(0)},{\mathcal{D}}_{-}^{(0)})$ with the higher/lower reward batch between ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. Next, we iteratively improve ${\mathcal{D}}_{+}$ through a swapping procedure. First, we tentatively swap a generation in ${\mathcal{D}}_{+}^{(0)}$ with one in ${\mathcal{D}}_{-}^{(0)}$ to form updated sets ${\mathcal{D}}_{+}^{{}^{\prime}(0)}$ and ${\mathcal{D}}_{-}^{{}^{\prime}(0)}$. Then, if ${\mathcal{D}}_{+}^{{}^{\prime}(0)}$ improves the reward over ${\mathcal{D}}_{+}^{(0)}$, then accept the swap and set $({\mathcal{D}}_{+}^{(1)},{\mathcal{D}}_{-}^{(1)})$ to $({\mathcal{D}}_{+}^{{}^{\prime}(0)},{\mathcal{D}}_{-}^{{}^{\prime}(0)})$. Otherwise, reject the swap and set $({\mathcal{D}}_{+}^{(1)},{\mathcal{D}}_{-}^{(1)})$ to $({\mathcal{D}}_{+}^{(0)},{\mathcal{D}}_{-}^{(0)})$. Repeating these steps, we obtain $({\mathcal{D}}_{+}^{(i)},{\mathcal{D}}_{-}^{(i)})$ for $i=0,1,\ldots$. When stopping conditions are met, we take the latest $({\mathcal{D}}_{+}^{(i)},{\mathcal{D}}_{-}^{(i)})$ as the final $({\mathcal{D}}_{+},{\mathcal{D}}_{-})$ pair. Each step of Algorithm 1 is guaranteed to improve or maintain ${\mathcal{D}}_{+}$’s reward. We maintain equal sizes for ${\mathcal{D}}_{1}$, ${\mathcal{D}}_{2}$, ${\mathcal{D}}_{+}$, and ${\mathcal{D}}_{-}$ and define the stopping condition as one complete pass over ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. When the loss function (to be discussed in Section 3.2) requires paired demonstrations, we use the same conditioning but different random seeds to generate ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. Subsequent swaps are also performed with same-conditioning pairs, ensuring that elements in ${\mathcal{D}}_{+}$ and ${\mathcal{D}}_{-}$ are paired to provide direct contrast.

### v2

@@ Abstract
We present Distributional RewArds for Generative OptimizatioN (DRAGON), a versatile framework for fine-tuning media generation models towards a desired outcome. Compared with traditional reinforcement learning with human feedback (RLHF) or pairwise preference approaches such as direct preference optimization (DPO), DRAGON is more flexible. It can optimize reward functions that evaluate either individual examples or distributions of them, making it compatible with a broad spectrum of instance-wise, instance-to-distribution, and distribution-to-distribution rewards. Leveraging this versatility, we construct novel reward functions by selecting an encoder and a set of reference examples to create an exemplar distribution. When cross-modal encoders such as CLAP are used, the reference may be of a different modality (e.g., text versus audio). Then, DRAGON gathers online and on-policy generations, scores them with the reward function to construct a positive demonstration set and a negative set, and leverages the contrast between the two finite sets to approximate distributional reward optimization. For evaluation, we fine-tune an audio-domain text-to-music diffusion model with 20 reward functions, including a custom music aesthetics model, CLAP score, Vendi diversity, and Fréchet audio distance (FAD). We further compare instance-wise (per-song) and full-dataset FAD settings while ablating multiple FAD encoders and reference sets. Over all 20 target rewards, DRAGON achieves an $81.45\%$ average win rate. Moreover, reward functions based on exemplar sets indeed enhance generations and are comparable to model-based rewards. With an appropriate exemplar set, DRAGON achieves a $60.95\%$ human-voted music quality win rate without training on human preference annotations. As such, DRAGON exhibits a new approach to designing and optimizing reward functions for improving human-perceived quality. Example generations can be found at https://ml-dragon.github.io/web.
@@ 1 Introduction
22footnotetext: Work done as an intern at Adobe Research, supported in part by the U.S. Army Research Laboratory and the U.S. Army Research Office under Grant W911NF2010219, Office of Naval Research, and NSF.
@@
Recent advances in diffusion models have transformed content generation across media domains, establishing new standards for generating high-quality images, video, and audio (LDM; VDM; AudioLDM; TANGO). While these models achieve impressive results through sophisticated training schemes, their optimization process typically focuses on metrics that may not align with downstream objectives or human preferences. This misalignment creates a fundamental challenge: how can we effectively steer these models toward desired output distributions or optimize them for specific performance metrics?
@@
A prominent approach to address this challenge has been fine-tuning using instance-level feedback. Methods such as reinforcement learning from human feedback (RLHF) (christiano2017deep; reinforce; PPO) and related methods like Direct Preference Optimization (DPO) (DPO) and Kahneman-Tversky Optimization (KTO) (KTO) leverage pre-trained reward models or large-scale, pairwise preference data collected offline to guide the optimization process. While effective, these approaches face several key challenges. Media like audio, music, and video are highly perceptual, making it hard and expensive to create reliable preference pairs (e.g. music (MusicRL), in-the-wild audio (BATON)). Research in language models considered implementing criteria-based reward signals to alleviate the preference data challenges (DeepSeekR1). However, similar approaches have been scarce for media creation, because its evaluation metrics often measure distributional properties like Fréchet embedding distance, diversity, and coverage, which existing reward optimization approaches like reinforcement learning could not handle. Furthermore, preference-based training methods like DPO (DPO) are constrained by the implicit reward functions hidden in their training data, making it difficult to adapt these approaches to new objectives or target distributions without collecting new preference data.
@@
To address these limitations, we introduce Distributional RewArds for Generative OptimizatioN (DRA-GON), a versatile framework for fine-tuning generative models towards a desired outcome or target distribution. DRAGON offers an alternative to existing reinforcement learning (RL) methods or pair-wise preference approaches for optimizing a broad spectrum of rewards, including instance-wise, instance-to-distribution, and distribution-to-distribution signals. As shown in Figure˜1(a), the key components of DRAGON are: 1) a pre-trained embedding extractor and a set of (possibly cross-modal) reference examples, 2) a reward-based scoring mechanism that creates positive and negative sets of on-policy online generations, and 3) an optimization process that leverages the contrast between the positive and negative finite sets to approximate distributional rewards. To our knowledge, DRAGON is the first practical algorithm that reliably optimizes our entire taxonomy of reward functions for generative models. We believe the ability to handle distribution-to-distribution rewards is particularly valuable since we can directly optimize generation quality metrics. Such reward functions are ubiquitous in generative model evaluation metrics (Fréchet embedding distance (FID), Kullback–Leibler (KL) divergence (KLD), Inception score (salimans2016improved)). Moreover, leveraging DRAGON’s unique versatility, we can construct new reward functions by simply collecting a set of ground-truth examples without human preference, drastically reducing the effort to construct reward signals.
@@
We demonstrate DRAGON’s effectiveness through comprehensive experiments with text-to-music diffusion transformers. Our evaluation incorporates multiple common music generation metrics: audio-text alignment (CLAP score) (clap; laionclap2023), Fréchet audio distance (FAD) (FAD) evaluated across diverse reference sets and embedding models, Vendi score (Vendi) for reference-free output diversity, and a custom-built human preference model for reference-free aesthetics scoring, totaling 20 reward functions. As shown in Figure˜1(b), DRAGON consistently improves over baselines across these diverse reward functions, achieving an average of 81.45% target reward win rate while generalizing the improvement across evaluation metrics. Via human listening tests, we show that DRAGON can achieve a 61.2% win rate in human-perceived music quality without training on human preference data.
@@
(a) Figure 1a: An overall diagram of DRAGON, a versatile on-policy learning framework for media creation models. DRAGON is compatible with diffusion-based generation and can optimize various types of distributional reward functions.
@@
(b) Figure 1b: DRAGON significantly improves a full suite of rewards. Each vertex considers a reward metric and reports the win rate of the DRAGON model optimized for the metric.
@@
  • We propose a human aesthetics preference model for AI-generated music and find that DRAGON can leverage it to improve human-perceived music quality with a small set of (1,676) human-rated clips.
@@ 2.1 Music Generation
Music generation has seen significant advances via auto-regressive token-based models (vaswani2017attention; zeghidour2021soundstream; MusicLM; AudioLM; MusicGen) and (latent) diffusion or flow-matching models (DITTO; noise2music; evans2024long; wu2024music; StableAudio; FlashAudio). Among auto-regressive approaches, MusicLM (MusicLM) and MusicGen (MusicGen) are notable, with the former extended with RL via MusicRL (MusicRL). Latent diffusion models like Stable Audio (evans2024long; StableAudio) show high-quality results and potential for ultra-fast generation (ConsistencyTTA; DITTO2; Presto). However, reward optimization for diffusion models has proven more challenging than for auto-regressive ones (Diffusion-DPO).
@@ 2.2 Diffusion Models
Diffusion models have emerged as a powerful paradigm able to generate high-quality and diverse samples (diffusion; DDPM) via iterative de-noising. Such models commonly consist of a de-noising network $f_{\theta}$ that inputs noisy input $x_{t}$, diffusion time step $t$, and condition $c$ (e.g. text), and can either be discrete-time (DDPM) or continuous-time score models (songscore; EDM). During training, given a demonstrative example $x_{0}$, we add Gaussian noise to form the noisy example $x_{t}$, and train $f_{\theta}$ to undo the noise injection. At inference time, generation begins with pure noise, from which $f_{\theta}$ is applied repeatedly to gradually denoise, forming a trajectory along the time step $t$ and noise level. Latent diffusion models (LDM), where the above diffusion process takes place in a latent embedding space, have been leveraged across nearly all perceptual generation domains, including creating images (LDM; EDM), video (VDM), speech (GradTTS; E2TTS), in-the-wild audio (AudioLDM; ConsistencyTTA), and music (noise2music; riffusion). Recent advances have shown that transformer-based architectures are advantageous, leading to diffusion transformers (DiT) (DiT).
@@ 2.3 Reward Optimization for Diffusion Models
Table 1: Comparison of methods that optimize diffusion models to reward functions or human preferences.
Category Example Methods for Diffusion Explicit Reward Distributional Reward
RLHF DDPO (DDPO),
DPOK (DPOK)
Preference Diffusion-DPO (Diffusion-DPO),
Optimization Diffusion-KTO (Diffusion-KTO)
DRAGON (ours) DRAGON
Diffusion models present unique challenges for reward-based optimization due to their iterative de-noising process. Existing work tackled this by formulating the diffusion process as a Markov Decision Process, with the reward signal assigned either exclusively to the final time step (DPOK) or to all time steps (DDPO). More recent works explored the alternative of implicit reward optimization via learning from contrastive demonstrations. Diffusion-DPO (Diffusion-DPO) and MaPO (MaPO) optimize rewards defined with explicit binary preferences between paired samples (e.g., media contents with the same captions). Diffusion-KTO (Diffusion-KTO) offers flexibility by being compatible with unpaired collections of preferred and non-preferred samples. Table˜1 presents a comparison between DRAGON and these existing methods. Concurrent to our work, TangoFlux (TANGOFlux) represents an application of a CLAP score reward in diffusion modeling, proposing a semi-on-policy approach for in-the-wild audio generation.
@@ 2.4 Human Feedback Datasets and Aesthetics Models
Human feedback datasets and aesthetics models play a crucial role in guiding the optimization of generative models towards human preferences and reward models. Datasets such as SAC (SAC), AVA (AVA), LAION-Aesthetics V2 (LAION-Aesthetics-Dataset), Pick-a-Pic (Pick-A-Pic), and RichHF-18K (liang2024rich) have been instrumental in improving the quality and alignment of generated images. These datasets are then used to train aesthetics models, which often map pre-extracted embeddings to a preference score. For audio/music generation, such an approach has been investigated with MusicRL (music) and BATON (in-the-wild audio), but is still rare and limited. In Section˜4.1 and Appendix˜B, we describe the construction of our own aesthetics dataset and reward model for music generation.
@@ 3 Distributional Reward Optimization For Diffusion Models
We propose DRAGON, a reward optimization framework for optimizing diffusion models with a wide variety of reward signals as shown in Figure˜1(a). We consider a distributional reward function $r_{\textrm{dist}}:{\mathcal{P}}\to{\mathbb{R}}$ that assigns a reward value to a distribution of generations, where $\mathcal{P}$} is the set of all such distributions. We allow $r_{\textrm{dist}}$ to be non-differentiable (e.g., human preference). The outputs of our generative model $f_{\theta}$ form a distribution ${\mathcal{D}}_{\theta}$. When $f_{\theta}$ is a conditional generator, ${\mathcal{D}}_{\theta}$ depends on the distribution of conditioning $\mathcal{C}$}, although we omit $\mathcal{C}$} for notation simplicity. Our goal is to fine-tune a pre-trained model via
@@
The formulation eq.˜1 includes traditional instance-level reward optimization widely studied in RLHF as a special case. Specifically, it is equivalent to instance-level optimization when the distributional reward $r_{\textrm{dist}}$ is the expectation of some instance-level reward $r_{\textrm{instance}}$, that is, $r_{\textrm{dist}}({\mathcal{D}}_{\theta})=\mathbb{E}_{X\sim{\mathcal{D}}_{\theta}}[r_{\textrm{instance}}(X)]$. However, unlike traditional RL, which is limited to instance-level rewards as it cannot distinguish between high and low quality generations without more granular feedback, DRAGON extends beyond this decomposable case.
To tackle such challenges and optimize $r_{\textrm{dist}}({\mathcal{D}}_{\theta})$, we construct a positive demonstrative distribution $\mathcal{D}$$_{+}$ such that $r_{\textrm{dist}}({\mathcal{D}}_{+})>r_{\textrm{dist}}({\mathcal{D}}_{\theta})$ and a negative demonstrative distribution $\mathcal{D}$$_{-}$. We then optimize model parameters $\theta$ to make ${\mathcal{D}}_{\theta}$ imitate $\mathcal{D}$$_{+}$ and repel $\mathcal{D}$$_{-}$. In the following sections, we address how to: 1) construct $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$, and 2) optimize $\theta$ to make ${\mathcal{D}}_{\theta}$ imitate $\mathcal{D}$$_{+}$ and repel $\mathcal{D}$$_{-}$. Please also find pseudocode in Appendix˜F.
@@ 3.1 On-Policy Construction of $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$
Existing work often assumes demonstrations $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$ to be provided in advance (offline and off-policy), with $\mathcal{D}$$_{+}$ known to achieve a higher reward than $\mathcal{D}$$_{-}$ (DPO; KTO; TANGO2). For instance-level rewards, this can be achieved by splitting a dataset into two halves at a reward threshold, a common RLHF approach. The limitations of off-policy learning necessitate a shift toward on-policy learning for several key reasons. First, for non-instance-level rewards such as FAD, the split becomes less straightforward. Second, large-scale offline data, required for effective off-policy learning, may be unavailable in practice. Third, on-policy optimization disentangles reward from dataset, providing more flexibility in reward choice. Finally, on-policy learning has demonstrated superior effectiveness and robustness because it enables real-time feedback and reduces data-policy mismatch. In the context of generative modeling, tajwar2024preference showed on-policy data helps language models learn from negative examples, and TANGOFlux showed that even a partially on-policy data collection pipeline improves diffusion models. Hence, we focus on an online and on-policy approach, although offline data can be optionally incorporated into DRAGON with minimal algorithmic changes.
@@
To construct on-policy distributions $\mathcal{D}$$_{+}$ and $\mathcal{D}$$_{-}$, we sample from ${\mathcal{D}}_{\theta}$ (which updates throughout training) online and query the distributional reward $r_{\textrm{dist}}$ on the fly. Specifically, before each training step, we collect two batches of observations from ${\mathcal{D}}_{\theta}$ by running full model inference with $f_{\theta}$ and denote them as ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. While the notations ${\mathcal{D}}_{1}$, ${\mathcal{D}}_{2}$, $\mathcal{D}$$_{+}$, and $\mathcal{D}$$_{-}$ technically represent distributions of generations, practical computations of $r_{\textrm{dist}}$ use sampled sets as approximations. Hence, with a slight abuse of notation, we also use these notations to denote the sampled demonstration sets. For the special case where the distributional reward $r_{\textrm{dist}}$ can be decomposed into some instance-level reward $r_{\textrm{instance}}$, that is, $r_{\textrm{dist}}({\mathcal{D}}_{\theta})=\mathbb{E}_{X\sim{\mathcal{D}}_{\theta}}[r_{\textrm{instance}}(X)]$, the contrastive demonstration sets $({\mathcal{D}}_{+},{\mathcal{D}}_{-})$ can be directly constructed by taking the better/worse halves of the union ${\mathcal{D}}_{1}\cup{\mathcal{D}}_{2}$. This split can be determined by protocols such as element-level pair-wise comparison (if ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$ consist of paired examples) or comparison with the batch median reward.
@@
Optimizing general non-decomposable rewards like $r_{\textrm{dist}}$ that evaluate distributions of generations is more delicate as we need to disentangle each element’s contribution. Exactly attributing each element’s contribution is combinatorially expensive and impractical to do at training time for each batch. To this end, we propose Algorithm˜1, a greedy algorithm. We initialize the positive/negative demonstration sets $({\mathcal{D}}_{+}^{(0)},{\mathcal{D}}_{-}^{(0)})$ with the higher/lower reward batch between ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$. Next, we iteratively improve $\mathcal{D}$$_{+}$ through a swapping procedure. First, we tentatively swap a generation in ${\mathcal{D}}_{+}^{(0)}$ with one in ${\mathcal{D}}_{-}^{(0)}$ to form updated sets ${\mathcal{D}}_{+}^{{}^{\prime}(0)}$ and ${\mathcal{D}}_{-}^{{}^{\prime}(0)}$. Then, if ${\mathcal{D}}_{+}^{{}^{\prime}(0)}$ improves the reward over ${\mathcal{D}}_{+}^{(0)}$, then accept the swap and set $({\mathcal{D}}_{+}^{(1)},{\mathcal{D}}_{-}^{(1)})$ to $({\mathcal{D}}_{+}^{{}^{\prime}(0)},{\mathcal{D}}_{-}^{{}^{\prime}(0)})$. Otherwise, reject the swap and set $({\mathcal{D}}_{+}^{(1)},{\mathcal{D}}_{-}^{(1)})$ to $({\mathcal{D}}_{+}^{(0)},{\mathcal{D}}_{-}^{(0)})$. Repeating these steps, we obtain $({\mathcal{D}}_{+}^{(i)},{\mathcal{D}}_{-}^{(i)})$ for $i=0,1,\ldots$. When stopping conditions are met, we take the latest $({\mathcal{D}}_{+}^{(i)},{\mathcal{D}}_{-}^{(i)})$ as the final $({\mathcal{D}}_{+},{\mathcal{D}}_{-})$ pair.